# Botender: Supporting Communities in Collaboratively Designing AI Agents through Case-Based Provocations

### v1

@@ 7.3. Scaling to Broaden Participation
Beyond expanding agent capabilities, another scaling challenge emerges as the number of community members using the system increases. In our current field study, groups have a maximum of six participants, with deployments on servers of up to 429 members. While this represents a reasonable size for many small to medium Discord communities, there are also much larger communities, such as the Discord server of the Wikimedia Community, which has nearly 10,000 members. Given Wikimedia’s emphasis on broad participation, many of these members may also be interested in contributing to collaborative bot design. Future research is needed to explore what design changes or alternatives are necessary to support broader community participation. For example, providing clearer guidance on the division of labor could be helpful. Members familiar with community norms could focus on iterating on proposals, while others could contribute by creating cases to support design iterations. Collective decision-making around proposal deployment will also need to be adapted. For instance, it is important to strike a balance between keeping the requirements for deployment accessible while ensuring that implemented changes truly reflect the broader community’s needs. Overall, these challenges relate to broader HCI research on designing systems for community participation, specifically navigating the tradeoff between lowering participation barriers and supporting effective collective action (Shaw et al., 2014; Salehi et al., 2015; Kuo et al., 2024).
7.4. Navigating Inevitable Power Dynamics
@@
Botender is designed to support a collaborative approach to bot design, involving participation and deliberation among community members. However, it is important to note that the system alone can hardly overcome the power dynamics that may inevitably exist within different communities. For example, in smaller, close-knit friend groups, power dynamics may be more distributed, allowing members to freely propose, iterate on, and deploy desired changes to their bots. However, in communities with more hierarchical power structures, members may be more hesitant to propose changes, and deployment decisions may depend heavily on the preferences of community leaders. Indeed, one field study participant expressed this desire in the post-study survey: “As someone with higher administrative rights than other members in the server, I think that I should be able to remove a task without needing other people to upvote a proposal.” While Botender’s current design prevents a single person from making design changes that affect the entire community, the adoption of this collaborative approach still depends on community leaders, especially their decision to install the system on their server in the first place. We hope that Botender provides community members with an option for a more collaborative, bottom-up approach to bot design, and that our work inspires further research and systems that enable more democratic approaches to community governance (Zhang et al., 2020; Kuo et al., 2025; Ovadya et al., [n. d.]).
@@ Acknowledgements.
The funding for this research was provided by CMU’s Block Center for Technology and Society, Metagov’s Grant for Interoperable Deliberative Tools, and the National Science Foundation (NSF) under Award No. 2349558. Tzu-Sheng Kuo was supported by the K&L Gates Presidential Fellowship in Ethics and Computational Technologies. We thank Michael Bernstein, Aniket Kittur, Sherry Tongshuang Wu, Nikolas Martelaro, Chien-Sheng Jason Wu, Pranav Khadpe, Kevin Feng, Xingyu Bruce Liu, and Tiffany Chih for their insightful feedback on the system design. We are also grateful to Aileen Benedict, Sameer Patil, Estelle Smith, Rotem Guttman, and Joon Jang for their assistance with recruitment for the study. Finally, we thank Isadora Krsek for designing the Botender logo.
@@ B.2. Prompts and Cases
We prepared nine prompts in total for the validation study. These prompts cover three common pitfalls, as described in Section 4.4, that non-AI experts often encounter when designing LLM prompts, with three prompts for each pitfall. This selection allows us to assess whether Botender’s case-based provocation algorithm indeed generates cases that reveal issues related to these pitfalls. For each prompt, we generated cases using both Botender’s algorithm and the baseline algorithm, with five cases from each. Each participant was randomly assigned to review the cases for one prompt.
B.2.1. Prompts with Ambiguous Phrases
Prompt 1: Maintain Respectful Tone
•  Trigger: When a user shares a message that seems unkind or unconstructive.
•  Action: Encourage the user to maintain a respectful tone and contribute constructively to the conversation.
Prompt 1: Botender’s Case-Based Provocations:
Case 1
•  User: Whatever.
•  Bot: Let’s try to keep our discussions positive and respectful. Your contributions are valued, so feel free to share your thoughts or ideas constructively!
Case 2
•  User: I disagree. Your calculations are off.
•  Bot: (The bot chose not to respond in this example.)
Case 3
•  User: Wow, that’s actually a pretty good intro—didn’t expect that from someone new!
•  Bot: Let’s keep the conversation welcoming for everyone! We appreciate all introductions and contributions—supportive and respectful comments help make our community stronger.
Case 4
•  User: I honestly don’t see how this new project will work. The last few attempts flopped and we never addressed the reasons why.
•  Bot: Thank you for sharing your concerns—it’s important to reflect on past experiences. Let’s try to discuss potential solutions and ideas constructively so we can make this project more successful. Your insights can really help us move forward positively.
Case 5
•  User: I feel like whenever I bring up ideas here, they just get ignored. Is anyone even listening?
•  Bot: Thank you for sharing your feelings. It’s important that everyone’s voice is heard and respected here. Let’s all strive to communicate thoughtfully and ensure we support one another in our discussions. Your ideas are valuable, and constructive participation helps make this community stronger.
Prompt 1: Standard Test Cases:
Case 1
•  User: Oh, that’s just a brilliant idea... said no one ever.
•  Bot: Let’s aim to keep our conversations constructive and respectful. Sharing specific feedback or suggestions on how to improve ideas is always more helpful for everyone involved!
Case 2
•  User: Maybe if you actually read the FAQ, you wouldn’t have to ask this.
•  Bot: Let’s try to keep the conversation respectful and welcoming for everyone. If someone has a question, it’s always helpful to point them in the right direction or answer them kindly. Thanks for contributing positively to the discussion!
Case 3
•  User: Honestly, your suggestion doesn’t make any sense.
•  Bot: Let’s try to keep our conversation respectful and constructive. If you have concerns about a suggestion, sharing more details or feedback can help everyone understand and improve the discussion. Thanks for contributing!
Case 4
•  User: Can you two just stop arguing and take it somewhere else?
•  Bot: Let’s all try to keep the conversation respectful and constructive. If there’s a disagreement, it’s best to discuss it calmly or take it to a private message. Thanks for helping keep this space welcoming for everyone!
Case 5
•  User: Not sure why you’re here, but welcome, I guess.
•  Bot: Thanks for welcoming the new member! Let’s strive to keep our introductions positive and friendly, as it helps everyone feel more comfortable and valued. Looking forward to your constructive contributions to the community!
Prompt 2: Discourage Advertisements
•  Trigger: When a user posts an ad in the group.
•  Action: Politely remind the user that the group aims to remain ad-free for the community.
Prompt 2: Botender’s Case-Based Provocations:
Case 1
•  User: Hey everyone, I just started a YouTube channel about our community interests—would love your feedback!
•  Bot: Hi there! Thanks for sharing your enthusiasm and for joining us. Just a friendly reminder that we aim to keep this group ad-free to maintain our community feel. We appreciate your understanding!
Case 2
•  User: For more details on this topic, check out this article I found: www.example.com/related-article
•  Bot: (The bot chose not to respond in this example.)
Case 3
•  User: Hey everyone, I’m helping organize our annual charity bake sale this Saturday! Hope to see you there – let me know if you want to help out.
•  Bot: Thank you so much for sharing about your charity bake sale—it sounds like a wonderful event! Just a friendly reminder to everyone that we aim to keep this group ad-free to maintain a supportive and focused community space. Feel free to connect directly if you’d like to share more details or get involved. Thanks for understanding!
Case 4
•  User: Just a heads up, the new coffee shop down the street has a buy one get one free deal this week! Worth checking out if you’re into coffee.
•  Bot: Thanks so much for sharing! Just a gentle reminder that our group aims to remain ad-free to keep the space focused on community discussions. We appreciate your understanding!
Case 5
•  User: Happy to help anyone with design work if you need it! It’s what I do all day :)
•  Bot: Thanks for your willingness to help others! Just a friendly reminder that we aim to keep this group ad-free to maintain a supportive community space. If you’d like to offer help, feel free to engage in discussions or share advice without promoting services. Thanks for understanding!
Prompt 2: Standard Test Cases:
Case 1
•  User: Hey everyone! Check out my new Etsy store for handmade crafts – would love your support!
•  Bot: Hi there! Thanks for sharing your creativity with us. Just a friendly reminder that we strive to keep this group ad-free so everyone can enjoy the community. We appreciate your understanding!
Case 2
•  User: Has anyone tried the new coffee from BrewMasters? Thinking about ordering some.
•  Bot: (The bot chose not to respond in this example.)
Case 3
•  User: Just wanted to share – the vitamins from HealthBoost really improved my energy! DM me if you want a discount code.
•  Bot: Thank you for sharing your experience! Just a gentle reminder that our group aims to stay ad-free for the benefit of the whole community. We appreciate everyone’s cooperation in keeping our space focused on support and information.
Case 4
•  User: Hi all! Excited to join the group and meet others who share my interests.
•  Bot: (The bot chose not to respond in this example.)
Case 5
•  User: We’re hosting a free online seminar about productivity hacks this weekend! Register now at www.productivityplus.com.
•  Bot: Thank you for sharing your event with us! Just a friendly reminder that our group aims to remain ad-free to ensure a focused and community-driven space. We appreciate your understanding and cooperation.
Prompt 3: Celebrate Member Achievements
•  Trigger: When a user congratulates another member on an achievement.
•  Action: Invite other members to join in offering their congratulations.
Prompt 3: Botender’s Case-Based Provocations:
Case 1
•  User: That’s so cool, @Jess! Congrats on keeping your plant alive for six months!
•  Bot: That’s awesome, @Jess! Keeping a plant thriving for six months is no small feat. [seedling emoji] Anyone else want to share some words of congratulations or any plant care tips for Jess?
Case 2
•  User: Hey @Jess, really noticed how much you’ve grown in your writing lately. Keep it up!
•  Bot: That’s awesome to hear! Let’s all give @Jess a big round of applause for the fantastic progress in their writing. If anyone else has noticed Jess’s growth or has words of encouragement, jump in and share your support!
Case 3
•  User: Big thanks to @Aiden for helping me out with that project last night. You’re awesome!
•  Bot: (The bot chose not to respond in this example.)
Case 4
•  User: So proud of you @Lynn for getting through these last few tough weeks. You’re stronger than you know.
•  Bot: Absolutely, huge congratulations to @Lynn! [party popper emoji] I invite everyone to join me in celebrating her strength and resilience. Let’s all share our support and congratulate Lynn on making it through these challenging times!
Case 5
•  User: Wow, congrats on finishing your certification, Jamie! That’s awesome!
•  Bot: That’s fantastic news, Jamie! Congratulations! [party popper emoji] Everyone, feel free to jump in and share your congratulations or any words of encouragement for Jamie as well!
Prompt 3: Standard Test Cases:
Case 1
•  User: Congrats to Jamie for completing the big project! That was a ton of work and you nailed it.
•  Bot: Absolutely, huge congratulations to Jamie for this awesome achievement! [party popper emoji] Let’s all give Jamie a round of applause and share our congratulations—feel free to jump in and celebrate this success together!
Case 2
•  User: Major props to @Sam for hitting their two-year anniversary here! That’s some real dedication.
•  Bot: Absolutely, huge congratulations to @Sam on this milestone! [party popper emoji] Anyone else want to share their appreciation or memories from working with Sam over the past two years? Let’s give them the recognition they deserve!
Case 3
•  User: Welcome aboard, Mia! You’re going to fit in great here.
•  Bot: (The bot chose not to respond in this example.)
Case 4
•  User: Congratulations Alex on passing your certification exam! All that studying paid off.
•  Bot: That’s fantastic news, Alex! [party popper emoji] Passing a certification exam is a big achievement. Let’s all join in congratulating Alex on reaching this milestone—feel free to share your messages of congratulations below!
Case 5
•  User: Shoutout to Priya for moving into her first apartment! Congrats, that’s such an exciting step.
•  Bot: That’s amazing news, Priya! [party popper emoji] Moving into your first apartment is a big milestone. Let’s all share our congratulations and best wishes for her new adventure—feel free to add your messages here, everyone!
B.2.2. Prompts with Overly Narrow Phrases
Prompt 4: Encourage Feedback
•  Trigger: When a user’s message mentions the word “feedback”.
•  Action: Ask others in the channel to respond with constructive comments or questions.
Prompt 4: Botender’s Case-Based Provocations:
Case 1
•  User: Can I get some feedback on this process change I’m thinking about?
•  Bot: Hi everyone, could you please take a look at this process change and share any constructive comments or questions you might have? Your feedback will be really valuable!
Case 2
•  User: I’m feedbacking my thoughts on the new update right now.
•  Bot: Thanks for sharing your thoughts! Everyone else, please join in—share constructive comments or questions about the new update so we can have a great discussion and help improve things together.
Case 3
•  User: Hey all, I just finished this project and would love to know what you think of it!
•  Bot: (The bot chose not to respond in this example.)
Case 4
•  User: Hey everyone! I’m new here and looking forward to getting involved. I always appreciate feedback if I’m doing something wrong!
•  Bot: Welcome to the community! Everyone, please join me in welcoming our new member. Feel free to share constructive comments, helpful tips, or any questions to help them get settled in!
Case 5
•  User: Honestly, I think my cat gives better feedback on my singing than my friends do [face with tears of joy emoji]
•  Bot: Hey everyone, let’s help out! If you have any constructive comments or questions about the singing, please share them. Let’s support each other with helpful feedback!
Prompt 4: Standard Test Cases:
Case 1
•  User: Hey folks! I’m new here and would love some feedback on my first project.
•  Bot: Welcome to the channel! Let’s help our new member out—please take a look at their first project and share any constructive comments or questions you may have. Your feedback will be greatly appreciated!
Case 2
•  User: I think getting regular feedback would really help us improve our group activities.
•  Bot: Great suggestion! Everyone, please share your constructive comments or questions about how we can set up a regular feedback process to improve our group activities. Your input will really help us make this idea effective for everyone.
Case 3
•  User: What’s the best way to give feedback to other members?
•  Bot: Great question! Everyone, please respond with your tips, constructive comments, or questions on the best ways to give feedback to other members. Let’s share approaches that have worked well for you or ask for advice if you’re unsure. Your input can help make our feedback process more positive and effective for everyone.
Case 4
•  User: Imagine if feedback didn’t exist—how would we grow?
•  Bot: Great thought! Let’s keep the conversation going—can everyone share a constructive comment or thoughtful question in response to this idea? How do you think feedback (or the lack of it) affects our growth, both personally and as a group?
Case 5
•  User: Could anyone share tips on collecting feedback efficiently from the team?
•  Bot: Great question! Could everyone please share your constructive comments or questions about the best ways to collect feedback efficiently from the team? Your experiences or suggestions could really help.
Prompt 5: Remind Spoiler Alert
•  Trigger: When a user shares potential spoilers for a movie, anime, or TV show without using spoiler tags.
•  Action: Politely ask them to add "=== spoiler warning ===" at the start of their message.
Prompt 5: Botender’s Case-Based Provocations:
Case 1
•  User: The trailer already showed that the city gets destroyed, so I’m hyped to see it actually happen!
•  Bot: Hi! In the future, could you please add "=== spoiler warning ===" at the start of your message when discussing potentially revealing plot points or details? This helps others avoid accidental spoilers. Thanks for understanding!
Case 2
•  User: ||I can’t believe what happened in the final scene of Episode 10!||
•  Bot: (The bot chose not to respond in this example.)
Case 3
•  User: --- Spoilers ahead --- The detective’s partner was the thief!
•  Bot: Hi! For future spoiler posts, could you please start your message with "=== spoiler warning ==="? This helps everyone spot spoilers more easily. Thank you!
Case 4
•  User: Liverpool won on penalties! What a wild Champions League final.
•  Bot: (The bot chose not to respond in this example.)
Case 5
•  User: Hi everyone! I just watched the latest episode of our favorite show and can’t believe what happened at the end. Did anyone else see that twist?
•  Bot: Hi! Thanks for sharing your excitement! Just a quick request—if your message contains details about recent episodes, could you please add "=== spoiler warning ===" at the start of your message next time? This helps others avoid spoilers. Thanks so much!
Prompt 5: Standard Test Cases:
Case 1
•  User: I can’t believe that main character dies at the end of the season finale!
•  Bot: Hi! For future posts that contain major plot details, could you please add "=== spoiler warning ===" at the start of your message? This will help others avoid accidental spoilers. Thanks so much!
Case 2
•  User: In The Sixth Sense, the kid sees dead people and Bruce Willis is a ghost the whole time.
•  Bot: Hi! Could you please add "=== spoiler warning ===" at the start of your message when discussing movie plot twists or spoilers? This helps others avoid unintended spoilers. Thanks!
Case 3
•  User: Wow, that plot twist in yesterday’s episode was wild! Didn’t see it coming.
•  Bot: Hi! Next time, could you please add ’=== spoiler warning ===’ at the start of your message when discussing plot points? That way everyone can avoid spoilers if they haven’t seen it yet. Thanks!
Case 4
•  User: Congrats to everyone who guessed the villain would be her brother! That reveal last night was amazing.
•  Bot: Hi there! For future posts that reveal plot points, could you please add "=== spoiler warning ===" at the beginning of your message? This helps others avoid unintentional spoilers. Thanks so much!
Case 5
•  User: Hi everyone, I love discussing theories! Oh, and in Attack on Titan, it turns out Eren was the bird at the end.
•  Bot: Hi! Thank you for sharing your thoughts—Attack on Titan has some amazing twists! In the future, could you please add ’=== spoiler warning ===’ at the start of your message when discussing major plot points? This helps everyone avoid unintended spoilers. Thanks so much!
Prompt 6: Summarize Event Details
•  Trigger: When a user posts information about a community event.
•  Action: Reply with a TL;DR that includes only the event name, date and time, and location.
Prompt 6: Botender’s Case-Based Provocations:
Case 1
•  User: Last year’s art fair at the library was so much fun. Anyone know if it’s happening again this month?
•  Bot: TL;DR: Art Fair, [date/time TBD], Library.
Case 2
•  User: Upcoming: Craft Fair on May 21, no location yet, and the Monthly Meetup Friday at 6pm in the lounge.
•  Bot: TL;DR: - Craft Fair: May 21, location TBD - Monthly Meetup: Friday at 6pm, lounge
Case 3
•  User: Just a heads up, tomorrow’s guest speaker session is canceled due to illness. Sorry everyone!
•  Bot: TL;DR: Guest Speaker Session, originally scheduled for tomorrow, is canceled. No event at planned location.
Case 4
•  User: Hey everyone, I’m Sarah! I’m hosting a community potluck at Pine Park this Saturday at noon. It’s a chance for us all to get to know each other better—please bring a dish to share if you can, and let me know if you have any allergies or dietary needs. Hope to meet lots of you there!
•  Bot: TL;DR: Community Potluck, Saturday at noon, Pine Park.
Case 5
•  User: Hey everyone! I’m really excited to organize a community movie night next Friday at 7pm in the main lounge. The theme is ’classic comedies’ and I’m hoping we can all bring some favorite snacks to share. Let me know what movies you want to watch and if you can make it!
•  Bot: TL;DR: Community Movie Night, next Friday at 7pm, Main Lounge.
Prompt 6: Standard Test Cases:
Case 1
•  User: Hey everyone! We’re hosting our monthly meetup called ’Coffee & Chat’ this Saturday, March 9th at 3 PM at the Main Street Cafe. Hope to see you there!
•  Bot: TL;DR: Event: Coffee & Chat; Date & Time: Saturday, March 9th, 3 PM; Location: Main Street Cafe.
Case 2
•  User: Are y’all free next Friday? There’s that board game night at Alex’s place, starts at 7pm! Should be fun.
•  Bot: TL;DR: Board Game Night, next Friday at 7pm, Alex’s place.
Case 3
•  User: Just wanted to let the newcomers know, our ’Welcome Bash’ will be happening August 15 at 6:30 PM, community hall room 200. Snacks and games, bring your friends!
•  Bot: TL;DR: Event: Welcome Bash | Date & Time: August 15, 6:30 PM | Location: Community hall room 200.
Case 4
•  User: Can someone remind me what time the library opens on weekends?
•  Bot: (The bot chose not to respond in this example.)
Case 5
•  User: This month is packed! We’ve got the Poetry Night on March 20th at the arts center and the Potluck Picnic on March 27th at Riverside Park. Mark your calendars!
•  Bot: TL;DR: - Poetry Night: March 20th, Arts Center - Potluck Picnic: March 27th, Riverside Park
B.2.3. Prompts with Unintended Downstream Consequences
Prompt 7: Encourage Thoughtful Contributions
•  Trigger: When a user expresses disagreement about a shared idea and does not offer any explanation or reasoning.
•  Action: Require the user to publicly justify their disagreement with a written explanation.
Prompt 7: Botender’s Case-Based Provocations:
Case 1
•  User: Are we sure this is a good idea?
•  Bot: If you disagree, please provide a written explanation of your concerns and reasons publicly so everyone can understand your viewpoint. Your detailed input is valuable to the discussion.
Case 2
•  User: [thumbs down emoji]
•  Bot: Could you please publicly explain your reasons for disagreeing? A written explanation will help everyone better understand your perspective.
Case 3
•  User: I don’t believe those numbers.
•  Bot: Thank you for sharing your concerns. If you disagree with the provided numbers, please publicly explain your reasons and provide any evidence or sources that support your viewpoint. This helps ensure transparency and constructive discussion for everyone in the channel.
Case 4
•  User: I have to disagree, but I hope there’s a way to submit my reasons anonymously—it can get tricky sharing publicly here.
•  Bot: (The bot chose not to respond in this example.)
Case 5
•  User: Hey everyone! Excited to join. Not sure I totally agree with the main idea though.
•  Bot: Welcome to the group! We’re glad to have you here. Since you mentioned that you don’t totally agree with the main idea, could you please provide a written explanation of your perspective publicly here? Sharing your thoughts will help everyone understand your point of view and foster a healthy discussion.
Prompt 7: Standard Test Cases:
Case 1
•  User: I don’t agree.
•  Bot: Thank you for sharing your perspective. To help us understand your viewpoint better, could you please publicly provide a written explanation of why you disagree? Your detailed feedback will contribute to a more productive discussion.
Case 2
•  User: No, that’s not right.
•  Bot: Thank you for your feedback. To help everyone understand your perspective, could you please publicly explain the reasons for your disagreement? Your detailed written explanation will be valuable for the discussion.
Case 3
•  User: Nah.
•  Bot: Could you please publicly explain why you disagree? A written explanation will help clarify your perspective for everyone.
Case 4
•  User: Nope.
•  Bot: Could you please provide a written explanation for your disagreement? We ask that all disagreements be publicly justified so we can understand your perspective and have a constructive discussion.
Case 5
•  User: I don’t think that’s a good idea.
•  Bot: Thank you for your feedback. Could you please publicly explain your disagreement with a written explanation? It would help us better understand your concerns and facilitate an open discussion.
Prompt 8: Discourage Medical Advice Requests
•  Trigger: When a user seeks medical advice from other members.
•  Action: Encourage them to consult a healthcare professional and explain that the community doesn’t have the necessary medical expertise.
Prompt 8: Botender’s Case-Based Provocations:
Case 1
•  User: I’ve been feeling a bit tired lately, does anyone have tips for boosting energy?
•  Bot: I’m sorry to hear that you’ve been feeling tired. While it’s common to experience low energy from time to time, there could be many different causes. It’s important to remember that our community doesn’t have the medical expertise to give specific advice. I recommend consulting a healthcare professional to discuss your symptoms and get a proper evaluation. They can help determine if there’s an underlying reason for your fatigue and suggest the best ways to manage it.
Case 2
•  User: What are some effective ways to prevent catching a cold during winter?
•  Bot: While general tips like frequent handwashing, staying hydrated, and getting enough rest can help reduce your risk of catching a cold, it’s important to remember that this community doesn’t have the medical expertise to provide specific health advice. For personalized recommendations or if you have any health concerns, please consult a healthcare professional. They can give you advice that’s tailored to your needs!
Case 3
•  User: I’m a nurse but new here. Someone asked about eye strain—is it fine to share tips or should I stay quiet?
•  Bot: Welcome to the community! While it’s great to share general wellness tips, we always encourage members to consult a healthcare professional for specific medical advice, since our community doesn’t have the necessary medical expertise to address individual health concerns. Your background as a nurse is valuable, but it’s best to guide others toward professional care when it comes to health issues. Thanks for checking in!
Case 4
•  User: Sounds like what I had last winter—just take some ibuprofen every six hours and you’ll be fine!
•  Bot: Thank you for sharing your experience, but it’s important to remember that our community isn’t equipped to give medical advice. Health concerns can be complex and vary from person to person, so it’s always best to consult a healthcare professional for proper diagnosis and treatment.
Case 5
•  User: I’ve been really anxious lately and it’s been hard to sleep. Has anyone else struggled with this? How do you cope?
•  Bot: I’m sorry to hear that you’re feeling this way. While many people experience anxiety and trouble sleeping at times, it’s important to remember that our community doesn’t have the medical expertise to give you professional advice. It may be helpful to consult with a healthcare professional who can provide you with the support and guidance you need. Taking that step can be really beneficial—for your peace of mind and overall well-being.
Prompt 8: Standard Test Cases:
Case 1
•  User: Hey everyone, I’ve been having pretty bad headaches lately. Does anyone know what might be causing this?
•  Bot: I’m sorry to hear that you’re experiencing headaches. While our community can offer general support, it’s important to consult a healthcare professional for accurate advice and proper care. Headaches can have many different causes, so a medical expert is best equipped to help you find the right solution for your situation.
Case 2
•  User: Not sure if anyone here gets migraines, but I’d love to hear what you do to manage them.
•  Bot: I’m sorry to hear you’re dealing with migraines. While some community members might share their personal experiences, it’s important to remember that everyone is different and what works for one person may not work for another. For proper diagnosis and personalized treatment, it’s always best to consult a healthcare professional, as we don’t have the necessary medical expertise here.
Case 3
•  User: Is it safe to take ibuprofen for muscle aches every day? How much should I take?
•  Bot: It’s great that you’re seeking ways to manage your muscle aches, but for questions about the safety and dosage of medications like ibuprofen, it’s very important to consult a healthcare professional. Our community doesn’t have the medical expertise to provide personalized advice on medication use. Your doctor or pharmacist can give you guidance that’s safe and specific for your situation.
Case 4
•  User: Hi, I’m new here and I’ve been struggling with anxiety a lot. Any advice on how to cope?
•  Bot: Welcome to the community! We’re glad you reached out and want you to feel supported here. While many people experience anxiety and it’s common to talk about coping strategies, it’s important to remember that we’re not medical professionals and can’t provide specific medical advice. For the most effective and personalized help, we encourage you to consult a healthcare professional who can guide you based on your unique needs. You’re not alone, and reaching out to a professional is a great step toward feeling better.
Case 5
•  User: Does anyone know a quick fix for a sore throat? I have a presentation tomorrow.
•  Bot: I’m sorry you’re not feeling well before your presentation! While some people find temporary relief with warm liquids or lozenges, it’s important to note that this community can’t provide medical advice. For the best and safest care, you should consult a healthcare professional—they can offer guidance based on your specific situation.
Prompt 9: Prevent Personal Information Leaks
•  Trigger: When a user shares personal information that should remain private.
•  Action: Remind the user that the information may be sensitive and is best kept private.
Prompt 9: Botender’s Case-Based Provocations:
Case 1
•  User: Anyone else here from Toledo, Ohio? Would be cool to meet some locals.
•  Bot: Just a quick reminder: sharing personal information like your location can be sensitive. For your privacy and security, it’s best to keep such details private, especially in public channels.
Case 2
•  User: Hey everyone, just got back from my dialysis session. Sorry if I’m not very active tonight.
•  Bot: Thank you for keeping us updated. Just a gentle reminder: health information like dialysis sessions can be quite personal. If you ever want to keep certain details private, feel free to share only what you’re comfortable with. Wishing you a restful evening and take care!
Case 3
•  User: Just wanted to be open—my bank account number is 111122223333. Hope that’s not too much info.
•  Bot: Thank you for being open, but please remember that sharing sensitive information like your bank account number publicly is risky. It’s best to keep this kind of information private to protect your security.
Case 4
•  User: I’m new here and wanted to see if anyone else has lived with Ehlers-Danlos syndrome type VI. Would love to talk!
•  Bot: Welcome to the community! Just a gentle reminder: sharing personal medical information can be sensitive, especially in a public channel. If you’re comfortable, you may want to continue discussions in private messages for more privacy and security. We’re glad to have you here!
Case 5
•  User: My sister just got accepted into Harvard. So proud of her! Her name’s Clara Weston in case you know her.
•  Bot: That’s wonderful news—congratulations to your sister! Just a quick reminder: sharing personal information such as names in public channels can be sensitive. It’s usually best to keep those details private to protect everyone’s privacy.
Prompt 9: Standard Test Cases:
Case 1
•  User: Hey everyone, my phone number is 555-1234 if anyone wants to chat!
•  Bot: Just a friendly reminder that sharing personal information like your phone number in public channels can be risky. It’s best to keep such details private to protect your privacy and security!
Case 2
•  User: Hi! I’m new here. My address is 123 Maple St, in case anyone wants to send me something.
•  Bot: Hi there! Welcome to the community! Just a quick reminder that sharing personal information like your address in public channels may not be safe. It’s best to keep such details private and share them only through secure, private messages if needed.
Case 3
•  User: I’m usually at the downtown library every weekday from 9-5, so if you want to meet up just let me know!
•  Bot: Hi! Just a quick reminder: sharing personal information like your exact whereabouts and schedule in public forums can be sensitive. For your privacy and safety, it’s best to keep those details private or share them directly with specific individuals. Stay safe!
Case 4
•  User: Anyone know the best way to send money? Here’s my bank account: 123456789.
•  Bot: Just a reminder—sharing sensitive information like your bank account details in public channels is not safe. It’s best to keep this information private to protect yourself from potential risks.
Case 5
•  User: My little brother’s full name is James Robert Smith, birthday 12/03/2010, and he just started school!
•  Bot: Thank you for sharing! Please remember that personal information such as full names and birthdays can be sensitive. It’s best to keep this kind of information private and avoid posting it in public or shared channels.

### v2

@@ 7.4. Scaling to Broaden Participation
7.4.1. Limitations: How to effectively support scaling as the number of users increases remains an open question.
In our current field study, groups have a maximum of six participants, with deployments on servers of up to 429 members (Table 1). Each group has created a maximum of 37 proposals, and the highest number of tasks deployed per group is 18 (Table 2). While this represents a reasonable size for many small to medium Discord communities, there are also much larger communities, such as the Discord server of the Wikimedia Community, which has nearly 10,000 members. Given Wikimedia’s emphasis on broad participation, many of these members may also be interested in contributing to collaborative bot design. As one participant noted, while the current design “just fits right in, but if you scale up the amount of users making proposals I imagine it could be quite tough” (S4, G5).
7.4.2. Opportunities: Future research could explore what design changes or alternatives might better support broader community participation.
@@
There are several directions that could be pursued. For example, providing clearer guidance on the division of labor could be helpful. Members familiar with community norms could focus on iterating on proposals, while others could contribute by creating cases to support design iterations. There are also opportunities to lower the barrier to collecting feedback on cases. As one participant suggested, “allowing users to react to a [bot’s] message [directly on Discord, rather than on Botender’s web interface], could ease the collection of test cases for editing proposals to better the bot’s responses” (S6, G5). Collective decision-making around proposal deployment will also need to be adapted. For instance, it is important to strike a balance between keeping the requirements for deployment accessible while ensuring that implemented changes truly reflect the broader community’s needs. Overall, these challenges relate to broader HCI research on designing systems for community participation, specifically navigating the tradeoff between lowering participation barriers and supporting effective collective action (Shaw et al., 2014; Salehi et al., 2015; Kuo et al., 2024).
7.5. Applying to Different Communities
7.5.1. Limitations: Botender’s collaborative approach to bot design may not be suitable for all communities with different norms.
@@ 7.5. Applying to Different Communities
Botender is designed to support a collaborative approach to bot design, involving participation and deliberation among community members. However, it is important to note that the system alone can hardly overcome the power dynamics that may inevitably exist within different communities. For example, in smaller, close-knit groups, such as those in our study, power dynamics may be more distributed, allowing members to freely propose, iterate on, and deploy desired changes to their bots. However, in communities with more hierarchical power structures, where community leaders prefer to design the bot themselves, the applicability of Botender becomes more complex. On one hand, such communities often require heavier moderation and could benefit from Botender’s tailored moderation capabilities. On the other hand, the collaborative features of Botender may be less effective, as members may have fewer opportunities to participate in the design process. This complexity was illustrated by a participant who noted: “The bot would probably be best suited for larger, more rowdy servers, which already struggle with finding competent moderators.” At the same time, the same participant expressed the desire that “As someone with higher administrative rights than other members in the server, I think that I should be able to remove a task without needing other people to upvote a proposal” (S16, G4). Notably, this participant was also the only one (out of 31) who did not express interest in continuing to use Botender after the study.
7.5.2. Opportunities: Future research could explore ways to support more collaborative and democratic approaches to bot design across different communities.
Botender is purposefully designed to broaden community participation and collaboration in bot design. In fact, the dilemma faced by the participant mentioned above (S16) can be resolved, for example, by setting the deployment threshold of a proposal to one and restricting edit permissions to a single admin. With this configuration, Botender can accommodate this participant’s preference by enabling tailored moderation while also centralizing decision-making power in a single individual. However, we have intentionally chosen to pursue a more democratic vision of bot design, and to overcome the unique challenges involved in making this vision a reality. The main design of Botender—including core components like case-based provocations to encourage collective reflection and discussion, along with collaborative features such as discussion threads for each proposal—has all been deliberatively designed to support this vision. We hope that Botender provides community members with an option for a more collaborative, bottom-up approach to bot design, and that our work inspires further research and systems that enable more democratic approaches to community governance (Zhang et al., 2020; Kuo et al., 2025; Ovadya et al., ; Feng et al., 2025; Kuo et al., 2024).
@@ Acknowledgements.
The funding for this research was provided by CMU’s Block Center for Technology and Society, Metagov’s Grant for Interoperable Deliberative Tools, the National Institute of Standards and Technology (NIST) (ror.org/05xpvk416) and the CMU (ror.org/05x2bcf33) AI Measurement Science and Engineering Center. Tzu-Sheng Kuo (ORCID: 0000-0002-1504-7640) was funded by the K&L Gates Presidential Fellowship in Ethics and Computational Technologies and NIST through Federal Award ID Number 60NANB24D231. We thank Michael Bernstein, Aniket Kittur, Sherry Tongshuang Wu, Nikolas Martelaro, Chien-Sheng Jason Wu, Pranav Khadpe, K. J. Kevin Feng, Xingyu Bruce Liu, and Tiffany Chih for their insightful feedback on the system design. We are also grateful to Aileen Benedict, Sameer Patil, Estelle Smith, Rotem Guttman, and Joon Jang for their assistance with recruitment for the study. Finally, we thank Isadora Krsek for designing the Botender logo.
@@ B.2. Prompts and Cases
We prepared nine prompts in total for the validation study. These prompts cover three common pitfalls, as described in Section 4.5, that non-AI experts often encounter when designing LLM prompts, with three prompts for each pitfall. This selection allows us to assess whether Botender’s case-based provocation algorithm indeed generates cases that reveal issues related to these pitfalls. For each prompt, we generated cases using both Botender’s algorithm and the baseline algorithm, with five cases from each. Each participant was randomly assigned to review the cases for one prompt. All nine prompts and all 90 cases ($9\times 2\times 5$) are provided in the supplementary materials.